# Algebraic Robustness Verification of Neural Networks

### v1

@@ B Proofs
B.2 Proof of Proposition˜3.2
@@
The first hidden layer $h_{1}\geq n$ produces a generic multivariate polynomial $P(\bm{x})$. The subsequent k1k-1 hidden layers of width 1 compute a univariate polynomial $G(t)$ of degree $d^{s-1}$, evaluated at $t=P(\bm{x})$. The final output layer is linear, so the decision boundary $B_{\bm{\theta}}(\bm{x})=f_{c}(\bm{x})-f_{c^{\prime}}(\bm{x})=0$ is given by:
@@
B.5 Proof of Theorem˜3.10
@@
$p_{F(\bm{x},\lambda)}(0)=p_{\left(f(\bm{x}),\nabla f(\bm{x})\right)}(0,\frac{\bm{\xi}-\bm{x}}{\lambda})\cdot|\det J_{\nabla f}(\bm{x})|=|\lambda|^{-n}\,p_{\left(f(\bm{x}),\nabla f(\bm{x})\right)}(0,\frac{\bm{\xi}-\bm{x}}{\lambda}).$

### v2

@@ A Verification Problem
B Numerical Solution via Homotopy Continuation
B.1 Formulation and the Gamma Trick
$H(\bm{z},t;\gamma):=\gamma(1-t)G(\bm{z})+tF(\bm{z}),$
B.2 Path Tracking via Predictor–Corrector Methods
B.3 Implementation and Post-Processing
C Experiments and Implementation Details
Benchmark configurations
Data generation
  • Set $\delta=m\cdot v$.
Model architecture and training
$\mathcal{L}=\mathcal{L}_{\text{CE}}+\mathcal{L}_{\text{margin}}$
where $f_{k}(x)$ denotes the logit for class $k$, and $m=0.01$ is the margin.
where $T=5,000$ is the total number of epochs.
Training Settings Results
6 2 0.2 100% 60% 0.0448(0.0063)0.0448\,(0.0063)
6 2 0.5 100% 60% 0.0546(0.0242)0.0546\,(0.0242)
6 3 0.2 100% 20% 5.8013(0.5377)5.8013\,(0.5377)
6 3 0.5 100% 40% 5.4401(0.2486)5.4401\,(0.2486)
10 2 0.2 100% 30% 0.0331(0.0063)0.0331\,(0.0063)
10 2 0.5 100% 50% 0.0335(0.0102)0.0335\,(0.0102)
10 3 0.2 100% 30% 4.6127(0.1721)4.6127\,(0.1721)
10 3 0.5 100% 50% 4.4646(0.2993)4.4646\,(0.2993)
Verification
For instances designated as unverifiable by construction, a sound verification method must always report falsification, since counterexamples are guaranteed to exist. Any verification claim on an unverifiable instance therefore constitutes a false positive and signals a soundness error in the verifier. In contrast, for clean instances, both verification and falsification are acceptable outcomes, depending on the true robustness of the model at the given input. This distinction allows us to separately evaluate verifier soundness on unverifiable instances and verifier precision on clean instances. Table˜3 summarizes the verification results across all 8 benchmark configurations. The verifier achieves perfect soundness on all unverifiable instances, correctly identifying all planted counterexamples. On clean instances, the verifier successfully falsifies a significant fraction, demonstrating its ability to detect non-robustness even without planted counterexamples. The average verification time per instance is also reported, showing efficient performance across configurations.
D ED Degree of the Decision Boundary of Polynomial Networks
D.1 Proof of Proposition˜3.2
@@ D ED Degree of the Decision Boundary of Polynomial Networks
The first hidden layer $h_{1}\geq n$ produces a generic multivariate polynomial $P(\bm{x})$. The subsequent s1s-1 hidden layers of width 1 compute a univariate polynomial $G(t)$ of degree $d^{s-1}$, evaluated at $t=P(\bm{x})$. The final output layer is linear, so the decision boundary $B_{\bm{\theta}}(\bm{x})=f_{c}(\bm{x})-f_{c^{\prime}}(\bm{x})=0$ is given by:
@@
E Expected Real ED Degree of the Decision Boundary of Polynomial Networks
E.1 Proof of Theorem˜3.10
@@ E Expected Real ED Degree of the Decision Boundary of Polynomial Networks
$p_{F(\bm{x},\lambda)}(0)=|\lambda|^{-n}\,p_{\left(f(\bm{x}),\nabla f(\bm{x})\right)}(0,\frac{\bm{\xi}-\bm{x}}{\lambda}).$